Suppose that the force on a 1kg mass on the surface of the earth is 10N and that the radius of the earth is about 600 km. Then the force on an 8kg mass placed at a point 2000 km from the center of the earth will be?


1
Expert's answer
2021-09-08T16:13:13-0400

"F =G\frac{mM}{R^2}"

"\text{for a mass of 1 kg on the ground:}"

"F_1 = 10N"

"R_1=600km=6*10^5m"

"F_1=G\frac{M}{(6*10^5)^2}"


"\text{for a mass of 8 kg at a distance of 2000 km}"

"\text{from the center of the earth:}"

"R_2= 2000km = 2*10^6m"

"F_2=G\frac{8M}{(2*10^6)^2}"

"F_2=8F_1\frac{(6*10^5)^2}{(2*10^6)^2}=7.2N"


"\text{Answer:}7.2N"